OpenStudy (anonymous):

What would the graph of y=x^2+1 look like? a. a straight like b. a parabola c. a dotted line d. none of the above

OpenStudy (anonymous):

Plug in values like I mentioned before and draw it out.

OpenStudy (anonymous):

I don't have a calc

OpenStudy (anonymous):

You don't need a calculator, you need to stop relying on the calculator.

OpenStudy (anonymous):

Then what do I have to do .-.

OpenStudy (anonymous):

Plug in 1 in for the x and see what the following output would be for y and do it for a decent amount of numbers, till you get the graph.

OpenStudy (anonymous):

Try what @iambatman suggested and if you do that and you still dont get it i can give you a hint

OpenStudy (anonymous):

(x,y) all you need put points

OpenStudy (anonymous):

I literally know nothing about this. ;-;

OpenStudy (anonymous):

You have to give it a try, before you give up at least.

OpenStudy (anonymous):

Idk how I would draw that

OpenStudy (anonymous):

ok just try when x = 0 what does y equal?

OpenStudy (anonymous):

IDK ;-;

OpenStudy (anonymous):

if you give it a try, just a random shot in the dark, iambatman or anyone else would know exactly what you know and dont know, and we can work from there. when you draw a graph, you can pick ANY number you want for either x or y. then you will get the coordinate pair that corresponds. you do this with three numbers you usually have a good idea of what is going on! trust iambatman because he is teaching you how to learn instead of this particular problem. y = x^2 +1

OpenStudy (anonymous):

plug 0 in for x! y = 0^2 +1. you should draw out a graph

OpenStudy (the_fizicx99):

o_o You don't need to plug in numbers, baman pls Look at the end-behaviors. It's quadratic, and the leading coefficient is positive so both sides will be point up. The +1 means it'll shift UP one interval on the "y" axis.

OpenStudy (anonymous):

y = 0^2 + 1 0^2 = 0 y = 0 + 1. do you understand that?

OpenStudy (the_fizicx99):

 Something like that :3

OpenStudy (anonymous):